# Robust Genomic Prediction and Heritability Estimation using Density Power Divergence

### v1

@@ 3.4 Real Data Analysis
We consider a real life breeding dataset generated by the International Maize and Wheat Improvement Center (CIMMYT) that contains 300 tropical maize lines in 10 blocks. Genotyping was conducted using 1135 SNP markers and the target phenotypic trait was grain yield assessed under severe drought stress and well-watered conditions. These two conditions are treated as two replicates in our analyses. Within the 300 tropical maize lines, 16 lines have missing values in some SNP markers. To avoid computational challenges due to missing values, these datapoints are excluded from our analysis presented in this paper. Thus all methods of genomic prediction, both existing and proposed, and associated heritability estimation are applied to the remaining 284 datapoints to predict their grain yields from 1135 SNP markers suitably adjusted from the block and the replicate effects. Since the underlying true values of the marker effects are unknown for a real dataset, in order to examine the performance of genomic predictions, we have computed the correlation coefficient $\widehat{\rho}$ and the MAD between the original and the predicted values of the target phenotypic trait, which is grain yield in our datset. The resulting values of $\widehat{\rho}$ and MAD, along with the estimated heritability measure $\widehat{H}_{p}$, are reported in Table 4 for different existing and proposed approaches.
@@ 4.1 Performance of the Proposed One-stage Method
Prediction accuracy measures, namely $\widehat{\rho}$ and MAD, obtained for different one-stage procedures are presented in Table 1 for artificial datsets generated under the first simulation setup; the results for pure data as well as two contaminated scenarios are presented. Examining these results, we observe that the predictive accuracy of the proposed MDPDE based approach improves ($\widehat{\rho}$ increases and MAD decrease) as $\alpha$ increases for both pure and contaminated situations; we get the best performances at $\alpha=1$ in all three cases.
Table 1: Prediction accuracy measures obtained from different methods under the one-stage approach with artificial datasets (first setup).
@@
RMLA 0.5216 26.0458 0.3754 46.452 0.2635 59.9679
RMLV 0.7182 5.0496 0.6008 3.962 0.3678 54.744
Rob-RMLA 0.5183 8.4920 0.4139 4.904 0.2898 83.6232
Rob-RMLV 0.7308 5.0485 0.6008 3.862 0.3678 53.0470
@@
$\alpha=0.1$ 0.8253 4.1311 0.7291 4.633 0.7180 13.2800
$\alpha=0.3$ 0.8255 4.1310 0.7292 4.632 0.7292 7.5980
$\alpha=0.5$ 0.8255 4.1301 0.7327 4.645 0.7341 6.9385
$\alpha=0.7$ 0.8257 4.0880 0.7395 4.604 0.7346 6.7637
$\alpha=1$ 0.8258 4.0430 0.7399 4.570 0.7352 6.4935
In the context of pure data, the proposed MDPDE based prediction method clearly outperforms other existing methods, namely both RMLA and RMLV and their robust counterparts. Although the performances of RMLV and Rob-RMLV are similar and better than RMLA and Rob-RMLA, they are yet significantly worse compared to our proposed MDPDE based approach which can achieve a high correlation of 0.8258 and a small MAD of 4.0430 at $\alpha=1$. Under contamination, the performances of classical method of RMLA and RMLV deteriorate drastically while those of their robust counterparts (Rob-RMLA and Rob-RMLV) remain somewhat stable in case of random contamination. In fact, Rob-MLVA achieves the smallest MAD value of 3.8624 among all methods under random contamination, making it a superior choice in terms of predictive accuracy measured by MAD. Nevertheless, the proposed MDPDE with $\alpha=1$ produces a competitive MAD value of 4.570 and a significantly higher correlation value (0.7399 compared to 0.6 of Rob-RMLV) and hence remains the best method in terms of correlation measure of accuracy. For block contaminated data, however, Rob-RMLV produces quite low correlation and high MAD values and the proposed MDPDE based procedure clearly emerges as the superior choice among all methods in terms of both predictive accuracy measures with improved performances as $\alpha$ increases. Comparing all set-ups and accuracy measures, the MDPDE based procedure remains most stable illustrating their significantly improved robustness properties over the existing procedures.
4.2 Performance of the Proposed Two-stage Method
The two-stage model analyses are conducted for the second experimental setup, and Table 2 presents the results of the predictive accuracy measures $\widehat{\rho}$ and the MAD for both pure and contaminated data. Based on the tabulated values, it is evident that the MDPDE method provides the best estimates among all the second stage methods and the effect of $\alpha$ is almost negligible on their performances although $\alpha=1$ gives marginally best results. Further, the results obtained by two-stage procedures are also significantly better than the one-stage results as intuitively expected; our proposed MDPDE based two-stage procedure can achieve an impressively high correlation of 0.9998, indicating strong correlation between the predicted and the original marker effects. Our method also leads to marker effect estimates having smaller deviations (MAD) from original values as compared with the existing Rob1 and Rob2 procedures. Further, the results from Table 2 also indicate that, when MDPDE is used in the second stage of the two-stage modeling, there is a significant reduction in MAD values and an increase in estimated correlation coefficients as compared with the existing for both block and response contaminated datasets. These observations suggest that the proposed MDPDE-based method exhibits even greater robustness compared to the only existing robust procedures.
Table 2: Prediction accuracy measures obtained from different methods under the two-stage approach with artificial datasets (second setup).
@@ 4.2 Performance of the Proposed Two-stage Method
$\alpha=0.1$ 0.9983 2.41 0.9914 2.679 0.9983 2.4074
$\alpha=0.3$ 0.9983 2.26 0.9915 2.678 0.9983 2.4069
$\alpha=0.5$ 0.9983 2.25 0.9923 2.678 0.9984 2.4066
$\alpha=0.7$ 0.9984 2.24 0.9924 2.675 0.9984 2.4064
$\alpha=1$ 0.9998 2.23 0.9925 2.679 0.9989 2.3062
@@
Figure 1: Plots of the accuracy measure $\widehat{\rho}$ obtained by the MDPDE-based two-stage genomic prediction (with $\alpha=1$) over different SNP marker-effect variance ($\sigma_{g}^{2}$) in the second experimental setup for two different $\sigma_{e}^{2}$.
However, it is to be noted that we have employed two different simulation setups for illustrating the accuracy of one-stage and two-stage modeling approaches. This is because, as we learned from extensive empirical explorations, performances of the proposed MDPDE based robust two-stage approach relies heavily on both error variance ($\sigma_{e}^{2}$) and (random) SNP-marker-effect variance ($\sigma_{e}^{2}$). To illustrate this dependence, in Figure 1, we have plotted the correlation measure ($\widehat{\rho}$) between the estimated and the original SNP marker-effects under the second experimental setup with different values of $\sigma_{e}^{2}$ and $\sigma_{e}^{2}$ for both pure and contaminated cases. It can be clearly seen from these figures (and many similar numerical explorations not presented here for brevity) that the accuracy of the proposed method increases as $\sigma_{g}^{2}$ increases and $\sigma_{e}^{2}$ increases for both pure and contaminated datasets, which is equivalent to say that the genotypic signal-to-noise ratio of the underlying model increases. Therefore, for datasets with a high genotypic effect variance compared to the noise variance, genomic prediction using the proposed MDPDE based two-stage modeling approach will be more accurate and robust. In other cases with lower SNP effect variance or higher noise variances, the use of robust one-stage procedure is expected to produce better results.
@@ 4.4 Accuracy and Robustness of the Heritability Estimator
In Table 3, we present the mean square deviation (MSD) of the heritability estimate ($\widehat{H}_{p}$), obtained by different existing and proposed methods in one-stage and two-stage approaches based on artificial pure and contaminated datasets in respective simulation set-ups. For one-stage analysis with the first experimental setup, the MSD of $\widehat{H}_{p}$ is minimized by the MDPDE method with $\alpha=0.3$ which is deemed to be the best choice for pure dataset as well as for noisy dataset with random contamination. But for the artificial dataset with incorporated block contamination, the desired minimum heritability achieved by the MDPDE method with $\alpha=1$ (which is $5.3731\times 10^{-10}$) in one-stage analysis. The proposed MDPDE-based two-stage analysis, however, produces significantly lower MSDs, indicating more accurate heritability estimation, for all $\alpha\geq 1$ in cases of both pure and contaminated datasets.
Table 3: The MSD (in $10^{-10}$) of heritability estimators obtained from different one-stage and two-stage procedures applied on pure and contaminated (cont.) experimental datasets.
Methods Pure Data Random Cont. Block Cont.
One-stage Procedures
RMLA 3.5235 5.378 8.3625
RMLV 98.1462 69.482 78.3400
ROB-RMLA 20.4525 9.7652 90.3215
ROB-RMLV 84.0401 59.3341 58.330
Proposed MDPDE based approach
$\alpha=0.1$ 1.7173 8.0626 7.3721
$\alpha=0.3$ 1.9353 2.5649 6.5677
$\alpha=0.5$ 2.9861 2.5407 8.4062
$\alpha=0.7$ 3.5678 5.1520 9.4702
$\alpha=1$ 4.3594 5.7077 5.3731
Two-stage Procedures
Rob1 19.7689 24.9496 25.8832
Rob2 60.2620 24.9373 25.0535
Proposed MDPDE based approach
$\alpha=0$ 2.9425 2.6456 2.9885
$\alpha=0.1$ 0.6341 0.2944 0.6324
$\alpha=0.3$ 0.4843 0.2982 0.7317
$\alpha=0.5$ 0.6657 0.3182 0.8079
$\alpha=0.7$ 0.9088 0.3677 0.8006
$\alpha=1$ 0.5741 0.4768 1.0044
@@ 4.5 Comparative Performances for the Maize Dataset
We have presented the correlation measure $\widehat{\rho}$ between the predicted and the true responses and their MAD obtained by different one-stage and two-stage methods in Table 4 along with the estimated values of the heritability measure ($\widehat{H}_{p}$). In case of this real data, we can see that the proposed MDPDE based one-stage modeling approach yields the best fit in terms of both correlation and MAD; the prediction performance further improves as $\alpha$ increases with $\alpha=1$ giving the highest correlation of 0.828 and the lowest MAD value 0.47, indicating minimal deviation between observed and predicted values. It’s worth noting that the two-stage approach have not produced better predictions for this dataset; this is because of the low marker effect variance compared to error variance as evident from extremely low estimated values of heritability given in Table 4. As per our observations presented in Section 4.3, one-stage modeling performs well when dealing with data exhibiting low variance in random SNP effects, which is the case in this real dataset. Consequently, one-stage model analysis with MDPDE proved to deliver desirable prediction accuracy measures.
Table 4: Prediction accuracy and heritability measures obtained by different one-stage and two-stage procedures applied on Maize dataset.
@@
RMLA 0.561 0.595 0.003770
RMLV 0.047 4.985 0.015600
ROB-RMLA 0.573 0.594 0.003760
ROB-RMLV 0.578 0.453 0.016400
@@
$\alpha=0.1$ 0.682 1.117 0.000666
$\alpha=0.3$ 0.684 1.116 0.001420
$\alpha=0.5$ 0.705 1.132 0.001550
$\alpha=0.7$ 0.710 1.132 0.001505
$\alpha=1$ 0.828 0.407 0.000706
@@
$\alpha=0.1$ 0.294 0.75 0.000196
$\alpha=0.3$ 0.282 1.88 0.000234
$\alpha=0.5$ 0.307 2.30 0.000266
$\alpha=0.7$ 0.311 2.26 0.000360
$\alpha=1$ 0.314 2.14 0.000384
@@ 5 Conclusions
In this paper, we have devised robust estimation techniques for genomic prediction based on an underlying linear mixed model. In pursuit of more robust estimation, our proposal also employs the popular MDPDE of the parameter of mixed models in two distinct modeling structure with one and two-stage approaches. We have compared our proposed methods to existing approaches using estimated marker effects in simulation experiments and predicted trait values in a real dataset. Remarkably, for both pure and contaminated data, the two-stage modeling structure consistently outperformed other approaches in terms of prediction accuracy whenever the random SNP-marker effects have greater variance compared to the random noise in the model. Our simulation study across various experimental setups demonstrates the sensitivity of results to the parameter $\alpha$ in the MDPDE based proposal with increasing value of $\alpha$ enhancing the accuracy and robustness of parameter and marker-effect estimation, ultimately yielding the more desirable predictive accuracy measure. Therefore, our comprehensive analysis establishes that, with a high correlation value and minimal deviation, the proposed MDPDE-based approaches with $\alpha=1$ proves to be the optimal choice for both one-stage and two-stage model structures offering more accurate and reliable gnomic prediction and heritability estimation compared to existing approaches. The promising results and insights derived from this research pave the way for exciting future prospects in the field of linear mixed models and genotypic-phenotypic effect estimation. Future studies can delve deeper into refining and improving the robustness of the proposed estimation methods. This could involve exploring and devising new techniques that adapt dynamically to data quality and contamination levels which would help to frame valuable contributions to genomics, genetics, and associated fields, ultimately advancing our comprehension of the intricate relationship between genetic factors and traits. We also hope to prepare an R-package for the proposed MDPDE-based procedures for genomic prediction and heritability estimation for dissemination among scientists and practitioners. They would then be able to apply our proposals easily in new experimental datasets to get more accurate and stable results, even in the presence of any possible noises and/or contamination, eventually leading to new insights and innovations.

### v2

@@ 3.4 Real Data Analysis
We consider a real life breeding dataset generated by the International Maize and Wheat Improvement Center (CIMMYT) that contains 300 tropical maize lines in 10 blocks. Genotyping was conducted using 1135 SNP markers and the target phenotypic trait was grain yield assessed under severe drought stress and well-watered conditions. These two conditions are treated as two replicates in our analyses.
Within the 300 tropical maize lines, 16 lines have missing values in some SNP markers. To avoid computational challenges due to missing values, these datapoints are excluded from our analysis presented in this paper. Thus all methods of genomic prediction, both existing and proposed, and associated heritability estimation are applied to the remaining 284 datapoints to predict their grain yields from 1135 SNP markers suitably adjusted from the block and the replicate effects.
In order to examine the performance of our proposed method for this real dataset, we have again computed the correlation coefficient $\widehat{\rho}$ and the MAD between the original and the predicted values of the target phenotypic trait, which is grain yield in our datset. The resulting values of $\widehat{\rho}$ and MAD, with 70-30 cross-validation as before, and the estimated heritability measure $\widehat{H}_{p}$, are reported in Table 4 for different existing and proposed approaches applied on the test data; corresponding results for the training data are given in the Appendix.
@@ 4.1 Performance of the Proposed One-stage Method
Prediction accuracy measures, namely $\widehat{\rho}$ and MAD, obtained for different one-stage procedures on the test data are presented in Table 1 for artificially generated pure data as well as two contaminated scenarios; similar results obtained for training data are given in Table 5 in the Appendix. Examining these results, we observe that the predictive accuracy of the proposed MDPDE based approach improves ($\widehat{\rho}$ increases significantly and MAD mostly decreases) as $\alpha$ increases for both pure and contaminated situations.
Table 1: Prediction accuracy measures obtained from different methods under the one-stage approach with artificial datasets (test data).
@@
RMLA 0.532 23.84 0.270 18.38 0.322 11.12
RMLV 0.587 16.77 0.387 16.69 0.470 18.38
Rob-RMLA 0.540 25.10 0.343 17.50 0.520 13.29
Rob-RMLV 0.712 7.23 0.430 8.18 0.550 6.48
@@
$\alpha=0.1$ 0.667 6.02 0.620 7.34 0.667 6.34
$\alpha=0.3$ 0.745 6.30 0.650 7.37 0.645 6.37
$\alpha=0.5$ 0.737 6.67 0.537 6.67 0.537 5.67
$\alpha=0.7$ 0.710 5.37 0.630 6.37 0.210 5.37
$\alpha=1$ 0.823 4.37 0.670 5.30 0.440 5.13
In the context of pure data, the proposed MDPDE based prediction method clearly outperforms other existing methods, namely both RMLA and RMLV and their robust counterparts, particularly in terms of MAD. In terms of $\widehat{\rho}$, the performances of RMLV and Rob-RMLV are similar and better than RMLA and Rob-RMLA, respectively, and also competitive with the MDPDE based methods with larger values of $\alpha$. But, they are significantly worse compared to our proposed MDPDE based approach which can achieve a significantly small MAD value between 4.37–6.67 at any $\alpha>0$.
Under contamination, the performances of classical method of RMLA and RMLV deteriorate drastically while those of their robust counterparts (Rob-RMLA and Rob-RMLV) remain somewhat stable. The proposed MDPDE based procedure with suitable $\alpha>0$ clearly emerges as the superior choice in terms of both predictive accuracy measures with improved performances as $\alpha$ increases.
Comparing all set-ups and accuracy measures, the MDPDE based procedure remains most stable, illustrating their significantly improved robustness properties over the existing procedures.
Table 2: Prediction accuracy measures obtained from different methods under the two-stage approach with artificial datasets (test data).
@@
$\alpha=0.1$ 0.820 4.02 0.760 4.14 0.747 4.34
$\alpha=0.3$ 0.845 4.20 0.750 3.67 0.742 4.37
$\alpha=0.5$ 0.837 3.67 0.737 3.39 0.733 4.67
$\alpha=0.7$ 0.810 2.37 0.820 3.08 0.758 3.37
$\alpha=1$ 0.860 2.37 0.832 2.93 0.800 3.34
4.2 Performance of the Proposed Two-stage Method
Table 2 presents the results of the predictive accuracy measures $\widehat{\rho}$ and the MAD for both pure and contaminated data with the two-stage modeling approach under test data; the same obtained for training data are given in Table 6 in the Appendix. Based on the tabulated values, it is evident that the MDPDE method with $\alpha\geq 0.5$ provides the best estimates among all the second stage methods and the effect of $\alpha$ is almost negligible on their performances although $\alpha=1$ gives marginally best results. Further, the results obtained by two-stage procedures are also significantly better than the one-stage results as intuitively expected; our proposed MDPDE based two-stage procedure can achieve a high correlation value of 0.86, indicating strong correlation between the predicted and the original phenotype traits (response). Our method also leads to predictions with smaller deviations (MAD) from their original values as compared with the existing Rob1 and Rob2 procedures. Further, the results from Table 2 (and Table 6) also indicate that, when MDPDE is used in the second stage of the two-stage modeling, there is a significant reduction in MAD values and an increase in estimated correlation coefficients as compared with the existing methods for both block and response contaminated datasets. These observations suggest that the proposed MDPDE-based method exhibits even greater robustness compared to the only existing robust procedures.
@@ 4.2 Performance of the Proposed Two-stage Method
Figure 1: Plots of the accuracy measure $\widehat{\rho}$ obtained by the MDPDE-based two-stage genomic prediction (with $\alpha=1$) over different SNP marker-effect variance ($\sigma_{g}^{2}$) in the second experimental setup for two different $\sigma_{e}^{2}$ (test data results).
However, it is to be noted that we have employed two different simulation setups for illustrating the accuracy of one-stage and two-stage modeling approaches. This is because, as we learned from extensive empirical explorations, performances of the proposed MDPDE based robust two-stage approach relies heavily on both error variance ($\sigma_{e}^{2}$) and (random) SNP-marker-effect variance ($\sigma_{g}^{2}$). To illustrate this dependence, in Figure 1, we have plotted the correlation measure ($\widehat{\rho}$) between the estimated and the original phenotype traits under the second experimental setup with different values of $\sigma_{e}^{2}$ and $\sigma_{g}^{2}$ for both pure and contaminated cases. It can be clearly seen from these figures (and many similar numerical explorations not presented here for brevity) that the accuracy of the proposed method increases as $\sigma_{g}^{2}$ increases and $\sigma_{e}^{2}$ decreases for both pure and contaminated datasets, which is equivalent to say that the genotypic signal-to-noise ratio of the underlying model increases. Therefore, for datasets with a high genotypic effect variance compared to the noise variance, genomic prediction using the proposed MDPDE based two-stage modeling approach will be more accurate and robust. In other cases with lower SNP effect variance or higher noise variances, the use of robust one-stage procedure is expected to produce better results.
Table 3: The average value and the MSD of heritability estimators obtained from different one-stage and two-stage procedures applied on pure and contaminated artificial (test) data.
Methods Pure Data Random Contamination Block Contamination
Average MSD Average MSD Average MSD
One-stage procedures
RMLA 0.4295 $7.3170\times 10^{-4}$ 0.4211 $7.7785\times 10^{-4}$ 0.4028 $8.8327\times 10^{-4}$
RMLV 0.5673 $1.7609\times 10^{-4}$ 0.5543 $2.1228\times 10^{-4}$ 0.5210 $3.2041\times 10^{-4}$
Rob-RMLA 0.6473 $2.7772\times 10^{-5}$ 0.5984 $1.0322\times 10^{-4}$ 0.5672 $1.7635\times 10^{-4}$
Rob-RMLV 0.6732 $7.1824\times 10^{-6}$ 0.6467 $2.8408\times 10^{-5}$ 0.6115 $7.8322\times 10^{-5}$
Proposed MDPDE based approach
$\alpha=0.1$ 0.7653 $4.264\times 10^{-5}$ 0.7276 $7.6176\times 10^{-6}$ 0.7112 $1.2544\times 10^{-6}$
$\alpha=0.3$ 0.7575 $3.3062\times 10^{-5}$ 0.7348 $1.211\times 10^{-5}$ 0.7154 $2.3716\times 10^{-6}$
$\alpha=0.5$ 0.7701 $4.914\times 10^{-5}$ 0.7365 $1.3322\times 10^{-5}$ 0.7261 $6.8121\times 10^{-6}$
$\alpha=0.7$ 0.8021 $1.0424\times 10^{-4}$ 0.7985 $9.7022\times 10^{-5}$ 0.7038 $1.444\times 10^{-7}$
$\alpha=1$ 0.7843 $7.1064\times 10^{-5}$ 0.7597 $3.564\times 10^{-5}$ 0.7387 $1.4976\times 10^{-5}$
Two-stage procedures
Rob1 0.7227 $5.1529\times 10^{-6}$ 0.7041 $1.681\times 10^{-5}$ 0.5768 $1.5178\times 10^{-4}$
Rob2 0.8022 $1.0444\times 10^{-4}$ 0.7836 $6.9889\times 10^{-5}$ 0.6363 $4.0576\times 10^{-5}$
Proposed MDPDE based approach
$\alpha=0.1$ 0.8190 $1.4161\times 10^{-4}$ 0.8030 $1.0609\times 10^{-4}$ 0.7260 $6.76\times 10^{-6}$
$\alpha=0.3$ 0.8230 $1.5129\times 10^{-4}$ 0.8050 $1.1025\times 10^{-4}$ 0.7332 $1.1022\times 10^{-5}$
$\alpha=0.5$ 0.8488 $2.2141\times 10^{-4}$ 0.8102 $1.2144\times 10^{-4}$ 0.7229 $5.2441\times 10^{-6}$
$\alpha=0.7$ 0.8670 $2.2788\times 10^{-4}$ 0.8254 $1.5725\times 10^{-4}$ 0.7481 $2.3136\times 10^{-5}$
$\alpha=1$ 0.8900 $3.61\times 10^{-4}$ 0.8543 $1.543\times 10^{-3}$ 0.7570 $3.249\times 10^{-5}$
@@ 4.4 Accuracy and Robustness of the Heritability Estimator
In Table 3, we present the average values and the mean square deviation (MSD) of the heritability estimate ($\widehat{H}_{p}$), obtained by different existing and proposed methods in one-stage and two-stage approaches based on artificial pure and contaminated datasets (test data). Corresponding results obtained under the training data are reported in Table 7 in the Appendix. In this respect, the proposed MDPDE based methods performs quite competitively to the existing robust methods, for both one-stage and two-stage approaches, under pure data as well as under randomly contaminated data. However, the MDPDE based methods show somewhat significant improvement in estimating the heritability over the existing methods under the block contamination scenario.
@@ 4.5 Comparative Performances for the Maize Dataset
We have presented the correlation measure $\widehat{\rho}$ between the predicted and the true responses and their MAD obtained by different one-stage and two-stage methods in Table 4 along with the estimated values of the heritability measure ($\widehat{H}_{p}$). In case of this real data, we can see that the proposed MDPDE based one-stage modeling approach yields the best fit in terms of both correlation and MAD; the prediction performance further improves as $\alpha$ increases with $\alpha=1$ giving the highest correlation of 0.671 and the lowest MAD value 0.418 in the test data, indicating minimal deviation between observed and predicted values.
It’s worth noting that the two-stage approach have not produced better predictions for this dataset; this is because of the low marker effect variance compared to error variance as evident from extremely low estimated values of heritability given in Table 4. As per our observations presented in Section 4.3, one-stage modeling performs well when dealing with data exhibiting low variance in random SNP effects, which is the case in this real dataset. Consequently, one-stage model analysis with MDPDE proved to deliver desirable prediction accuracy measures.
Table 4: Prediction accuracy and heritability measures obtained by different one-stage and two-stage procedures applied on Maize dataset (test data results).
@@
RMLA 0.404 0.658 0.006138
RMLV 0.117 0.748 0.014968
Rob-RMLA 0.416 0.557 0.003128
Rob-RMLV 0.421 0.421 0.016768
@@
$\alpha=0.1$ 0.525 1.106 0.003034
$\alpha=0.3$ 0.527 1.879 0.003788
$\alpha=0.5$ 0.548 1.943 0.003918
$\alpha=0.7$ 0.553 1.704 0.003873
$\alpha=1$ 0.671 0.418 0.003074
@@
$\alpha=0.1$ 0.390 1.697 0.00576
$\alpha=0.3$ 0.378 1.827 0.00580
$\alpha=0.5$ 0.303 1.247 0.00584
$\alpha=0.7$ 0.407 1.207 0.00593
$\alpha=1$ 0.410 1.087 0.00595
@@ 5 Conclusions
In this paper, we have devised robust estimation techniques for genomic prediction based on an underlying linear mixed model. In pursuit of more robust estimation, our proposal also employs the popular MDPDE of the parameter of mixed models in two distinct modeling structure with one and two-stage approaches. We have compared our proposed methods to existing approaches using estimated marker effects in simulation experiments and predicted trait values in a real dataset. Remarkably, for both pure and contaminated data, the two-stage modeling structure consistently outperformed other approaches in terms of prediction accuracy whenever the random SNP-marker effects have greater variance compared to the random noise in the model. Our simulation study across various experimental setups demonstrates the sensitivity of results to the parameter $\alpha$ in the MDPDE based proposal with increasing value of $\alpha$ enhancing the accuracy and robustness of parameter and marker-effect estimation, ultimately yielding the more desirable predictive accuracy measure. Therefore, our comprehensive analysis establishes that, with a high correlation value and minimal deviation, the proposed MDPDE-based approaches with $\alpha=1$ proves to be the optimal choice for both one-stage and two-stage model structures offering more accurate and reliable gnomic prediction and heritability estimation compared to existing approaches.
We would like to point out that the MDPDE indeed robustifies the maximum likelihood (ML) estimator and thus our approaches utilize a robust ML-type estimator of variance parameters to make genomic prediction under noisy data. Although this approach is seen to produce better prediction performances compared to the existing competitors, the ML-type variance estimators are known to yield biased variance estimates under linear mixed models while the restricted ML (REML) estimator yields an improved variance estimate. Unfortunately, however, when focusing on robustness against data contamination, there is still no method available in the literature for producing a robust REML-type variance estimators from possibly contaminated data under the linear mixed models. It would be a significantly important future research work to develop such a suitable robust REML-type estimation procedure under general linear mixed models and subsequently incorporate it within our proposed approaches to further improve the accuracy of robust genomic prediction.
The promising results and insights derived from this research pave the way for many other exciting future prospects in the field of linear mixed models and genotypic-phenotypic effect estimation. This could involve exploring and devising new techniques that adapt dynamically to data quality and contamination levels, e.g., data-driven algorithm for selecting optimum value of the tuning parameter $\alpha$. The basic heritability measure $\widehat{H}_{p}$ can also be improved further accounting for selection, along the line of, e.g., Schmidt et al. (2019) [29] and Feldmann et al. (2023) [11]. All these would eventually lead to comprehensive understanding of the intricate relationship between genetic factors and traits.
We also hope to prepare an R-package for the proposed MDPDE-based procedures for genomic prediction and heritability estimation for dissemination among scientists and practitioners. They would then be able to apply our proposals easily in new experimental datasets to get more accurate and stable results, even in the presence of any possible noises and/or contamination, eventually leading to new insights and innovations.
Appendix A Numerical Results obtained on the training datasets
We have presented all the numerical results obtained under the training data in Tables 58. The relative performances of different one-stage and two-stage methods are again the same as observed in the test data, although their individual performances are much improved for the training data than those seen under test data. This is intuitively expected as the models are fitted with the training data. However, most importantly, our proposed MDPDE based methods perform quite closely under both their training data as well as the test data.
Table 5: Prediction accuracy measures obtained from different methods under the one-stage approach with artificial datasets (training data).
Methods Pure Data Random Contamination Block Contamination
$\widehat{\rho}$ MAD $\widehat{\rho}$ MAD $\widehat{\rho}$ MAD
RMLA 0.366 37.77 0.238 35.70 0.235 52.87
RMLV 0.687 4.11 0.544 4.107 0.441 9.56
Rob-RMLA 0.410 7.50 0.588 6.77 0.472 13.64
Rob-RMLV 0.648 5.46 0.650 4.97 0.465 27.67
Proposed MDPDE based approach
$\alpha=0.1$ 0.767 4.34 0.723 4.48 0.623 8.02
$\alpha=0.3$ 0.745 4.37 0.766 4.36 0.666 6.38
$\alpha=0.5$ 0.737 5.67 0.734 4.67 0.534 6.07
$\alpha=0.7$ 0.710 4.37 0.710 4.30 0.730 6.04
$\alpha=1$ 0.813 4.06 0.774 4.04 0.745 5.37
Table 6: Prediction accuracy measures obtained from different methods under the two-stage approach with artificial datasets (training data).
Methods Pure Data Random Contamination Block Contamination
$\widehat{\rho}$ MAD $\widehat{\rho}$ MAD $\widehat{\rho}$ MAD
Rob1 0.756 5.43 0.788 4.07 0.772 3.64
Rob2 0.808 5.87 0.772 4.87 0.765 3.37
Proposed MDPDE based approach
$\alpha=0.1$ 0.837 4.34 0.803 4.04 0.799 3.22
$\alpha=0.3$ 0.865 4.37 0.820 3.48 0.796 3.38
$\alpha=0.5$ 0.877 3.67 0.805 3.67 0.802 3.27
$\alpha=0.7$ 0.840 3.77 0.830 3.37 0.814 3.33
$\alpha=1$ 0.884 3.07 0.833 3.30 0.823 3.18
Table 7: The average value and the MSD of heritability estimators obtained from different one- and two-stage procedures applied on pure and contaminated artificial (training) data.
Methods Pure Data Random Contamination Block Contamination
Average MSD Average MSD Average MSD
One-stage procedures
RMLA 0.4543 $6.0368\times 10^{-4}$ 0.3181 $1.4584\times 10^{-3}$ 0.3018 $1.5856\times 10^{-3}$
RMLV 0.5737 $1.5951\times 10^{-4}$ 0.5492 $2.274\times 10^{-4}$ 0.5175 $3.3306\times 10^{-4}$
Rob-RMLA 0.6345 $4.2902\times 10^{-5}$ 0.6190 $6.561\times 10^{-5}$ 0.5939 $1.1257\times 10^{-4}$
Rob-RMLV 0.6861 $1.9321\times 10^{-5}$ 0.6332 $4.4622\times 10^{-5}$ 0.6303 $4.858\times 10^{-5}$
Proposed MDPDE based approach
$\alpha=0.1$ 0.6927 $5.329\times 10^{-7}$ 0.6570 $1.849\times 10^{-5}$ 0.6373 $3.9312\times 10^{-5}$
$\alpha=0.3$ 0.7515 $2.6522\times 10^{-5}$ 0.7229 $5.2441\times 10^{-6}$ 0.6408 $3.5046\times 10^{-5}$
$\alpha=0.5$ 0.7080 $6.4\times 10^{-7}$ 0.6760 $5.76\times 10^{-6}$ 0.6529 $2.2184\times 10^{-5}$
$\alpha=0.7$ 0.7306 $9.3636\times 10^{-6}$ 0.7104 $1.0816\times 10^{-6}$ 0.6976 $5.76\times 10^{-8}$
$\alpha=1$ 0.6974 $5.76\times 10^{-8}$ 0.6633 $1.3468\times 10^{-5}$ 0.7038 $1.444\times 10^{-7}$
Two-stage procedures
Rob1 0.7796 $6.3361\times 10^{-5}$ 0.7538 $2.8944\times 10^{-5}$ 0.7630 $3.969\times 10^{-5}$
Rob2 0.7906 $8.2083\times 10^{-5}$ 0.7745 $5.5502\times 10^{-5}$ 0.7350 $1.225\times 10^{-5}$
Proposed MDPDE based approach
$\alpha=0.1$ 0.7658 $4.4329\times 10^{-5}$ 0.7492 $2.4206\times 10^{-5}$ 0.6850 $5.625\times 10^{-5}$
$\alpha=0.3$ 0.8100 $1.21\times 10^{-4}$ 0.8003 $1.006\times 10^{-4}$ 0.7280 $7.84\times 10^{-6}$
$\alpha=0.5$ 0.8281 $1.64\times 10^{-4}$ 0.8010 $1.02\times 10^{-4}$ 0.7910 $8.281\times 10^{-5}$
$\alpha=0.7$ 0.8435 $2.0592\times 10^{-4}$ 0.8218 $1.4835\times 10^{-4}$ 0.8030 $1.0609\times 10^{-4}$
$\alpha=1$ 0.8670 $2.7889\times 10^{-4}$ 0.8340 $1.7956\times 10^{-4}$ 0.8220 $1.4884\times 10^{-4}$
Table 8: Prediction accuracy and heritability measures obtained by different one-stage and two-stage procedures applied on Maize dataset (training data results).
One-stage Procedures
Method $\widehat{\rho}$ MAD $\widehat{H}_{p}$
RMLA 0.561 0.591 0.003771
RMLV 0.047 0.581 0.01640
Rob-RMLA 0.577 0.590 0.003561
Rob-RMLV 0.554 0.474 0.016401
Proposed MDPDE based approach
$\alpha=0.1$ 0.664 1.139 0.000666
$\alpha=0.3$ 0.6837 1.12 0.001422
$\alpha=0.5$ 0.689 1.172 0.001351
$\alpha=0.7$ 0.716 1.537 0.001306
$\alpha=1$ 0.82 0.453 0.000706
Two-stage Procedures
Method $\widehat{\rho}$ MAD $\widehat{H}_{p}$
Rob1 0.726 2.430 0.002310
Rob2 0.726 2.420 0.002050
Proposed MDPDE based approach
$\alpha=0.1$ 0.294 0.750 0.000196
$\alpha=0.3$ 0.282 1.880 0.000234
$\alpha=0.5$ 0.307 2.300 0.000266
$\alpha=0.7$ 0.311 2.260 0.000360
$\alpha=1$ 0.314 2.140 0.000384